OpenStudy (anonymous):

I don't know how to get the answer? Let alone do it:( Containing E(4, 3) and F(6, 1)

OpenStudy (anonymous):

find linear equation?

OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation in standard form.

OpenStudy (anonymous):

Do you know Point slope form?

OpenStudy (anonymous):

Yes, but I don't fully understand how to get the answer

OpenStudy (anonymous):

Show me the formula, pls!

OpenStudy (anonymous):

for point slope?

OpenStudy (anonymous):

Yes, that's the most essential form for linear equation!

OpenStudy (anonymous):

if I understand correctly it is y=1/2 * x/2 + something?

OpenStudy (anonymous):

the formula, with slope m!

OpenStudy (anonymous):

oh... um I don't know that one.

OpenStudy (anonymous):

Would you look it up for the formula ( without the number)

OpenStudy (anonymous):

look what up?

OpenStudy (anonymous):

Point slope form for linear from where assign your homework?

OpenStudy (anonymous):

I go to school online, and nothing was given about that. Just this equation as if they thought every student knew how to do it.

OpenStudy (anonymous):

Yes, everybody are supposed to memorize it! y2 - y1 = m ( x2 - x1 )

OpenStudy (anonymous):

So I would put the numbers in as the letters, and than get the answer that way?

OpenStudy (anonymous):

First you find slope m = .... Plug the coordinate in:

OpenStudy (anonymous):

how do I know which ones to use to find the slope?

OpenStudy (anonymous):

Slope m = ( y2 - y1) / ( x2 - x1) You need TWO POINTS to make a line

OpenStudy (anonymous):

So y2 would be 4 and y1 would be 6? and x2 would be 3 and x1 would be 1? or switch it around?

OpenStudy (anonymous):

E(4, 3) and F(6, 1) x1 = 4, y1 = 3, x2 = 6, y2 = 1

OpenStudy (anonymous):

x=2 / y=-2 ?

OpenStudy (anonymous):

is that the answer? or do I divide 2 and -2?

OpenStudy (anonymous):

I have 5spaces in the answer box...and I think I did something wrong

OpenStudy (anonymous):

Oh...so it would be y=-2 as the answer?

OpenStudy (anonymous):

x2-x1= 6-4=2

OpenStudy (anonymous):

m=-1?

OpenStudy (anonymous):

Yes, now plug into the linear equation y - y1 = m ( x - x1 )

OpenStudy (anonymous):

ok I think IM over thinking this....what would I plug in? would it be -2=m2? but I can only Y & X in the answer?

OpenStudy (anonymous):

You plug in ( 6,1) and m = -1

OpenStudy (anonymous):

im lost? im sorry...I hope im not aggravating you

OpenStudy (anonymous):

You really don't know point coordinate, don't you! x1 = 6, y1 = 1, m = -1 Now plug these into the formula: y - y1 = m ( x - x1 )

OpenStudy (anonymous):

I know the coordinates to the point of finding where the points go on the graph, but I never was showed how to find them.

OpenStudy (anonymous):

So. pls follow my instruction Point coordinate is x1 = 6, y1 = 1, m = -1 Now plug these into the formula: y - y1 = m ( x - x1 )

OpenStudy (anonymous):

x=4, y=3= 1, m=1?

OpenStudy (anonymous):

2=m-1

OpenStudy (anonymous):

Either point, it works, m -1

OpenStudy (anonymous):

y - y1 = y - 3 Can you do similarly with x

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (anonymous):

x - x1 = ...

OpenStudy (anonymous):

3-1=2

OpenStudy (anonymous):

x - 4

OpenStudy (anonymous):

subtract 3 from 4?? =-1

OpenStudy (anonymous):

y - 3 = - ( x - 4 )

OpenStudy (anonymous):

y -3 = -x + 4 ->x = -x + 4 + 3 x = -x + 7